OpenStudy (anonymous):

To convert degrees Fahrenheit (F) into degrees Celsius (C), use the formula C=5/9(F-32). Convert 15 °C to Fahrenheit.

OpenStudy (anonymous):

i say it is 59

OpenStudy (charlotte123):

@msumner u r right :D c = 15. So, the equation would be 15=5/9(F-32). Convert 15 °C to Fahrenheit. substituting 59 for F, we get 15=5/9(59-32) = 15 =5/9(27). 27 * 5/9 = 15, so the statment is true :D thus, 15 °C = 59 °F :D

OpenStudy (anonymous):

thanks

OpenStudy (charlotte123):

@msumner my pleasure :D have anymore questions? :)